In a time and motion study, it was found that a certain manual work can be finished at an average time of 40 minutes with a standard deviation of 8 minutes. A group of 16 students is given a special training and then found to average only 35 minutes. Can we conclude that the special training can speed up the work using 0.01 significance level?


1
Expert's answer
2021-03-31T02:01:48-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\geq40"

"H_1:\mu<40"

This corresponds to a left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.01," and the critical value for a left-tailed test is "z_c=-2.3263."

The rejection region for this left-tailed test is "R=\{z:z<-2.3263\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{X}-\mu}{\sigma/\sqrt{n}}=\dfrac{35-40}{8/\sqrt{16}}=-2.5"

Since it is observed that "z=-2.5<-2.3263=z_c," it is then concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is less than 40, at the 0.01 significance level.


Using the P-value approach: The p-value is "p=P(Z<-2.5)=0.0062," and sice "p=0.0062<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is less than 40, at the 0.01 significance level.

We can conclude that the special training speed up the work using "\alpha=0.01".